The manager of small candy shop operatinh in perfectly competitive market determines that his average cost for chocolates is given by 10-0.2Q +0.005Q^2 where Q is in pound per month. The market price of chocolates is expected to maintain at $8 per pound. How many pounds per month should he produce?
1
Expert's answer
2019-05-27T11:56:22-0400

In the long run P = ATC, so:

10 - 0.2Q + 0.005Q^2 = 8,

Q^2 - 40Q + 400 = 0,

(Q - 20)^2 = 0,

Q1 = 20, Q2 = 0.